What value resistor will discharge a 2.10 μF capacitor to 20.0% of its initial charge in 2.30 ms ?
Answer #1

the equation for discharging of capacitor in RC circuit is

Q= Qo e^- t/RC

(20/100) Q0 = Q0 e^-2.3* 10^-3 s/R(2.10* 10^-6)

-2.3* 10^-3 s/ R (2.1* 10^-6) = ln(0.2)

R = 680.5 ohm